OpenStudy (anonymous):

$4000 is invested in a bank account at an interest rate of 8% per year, compounded continuously. Meanwhile, $17000 is invested in a bank account at an interest rate of 3% compounded annually. To the nearest year, When will the two accounts have the same balance? After how many years?

OpenStudy (campbell_st):

well continuous compounding is \[A = P \times e^{rt}\] A = future value, P = principal r = interest rate as a decimal and t = time plug the values in

OpenStudy (anonymous):

im wasnt sure which one to use

OpenStudy (campbell_st):

the annual compound interest formula is \[A = P \times (1 + r)^t\] the letter have the same definitions... just plug the numbers in

OpenStudy (anonymous):

but one of them says comopunded annually

OpenStudy (campbell_st):

so then you need to equate both formulae and then solve for t. to find the time when the balances are the same..

OpenStudy (anonymous):

so for the first one i got 4000e^(.08*t)

OpenStudy (anonymous):

is that part right?

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

so the second equation is 17000(1.03)^t right?

OpenStudy (campbell_st):

that's correct... so the equation becomes \[4000e^{0.8t} = 17000(1.03)^t\] divide by 4000 \[e^{0.8t} = 4.25\times (1.03)^t\] now you need to take the log of both sides... so use base e logs or ln.. hope it helps

OpenStudy (anonymous):

oh ok i get it, its cuz when i divided by 4000 i put it under (1.03)^t as well

OpenStudy (campbell_st):

you well need to know about 1. log laws for multiplication 2. log laws for powers 3. log laws for the log of the base...

OpenStudy (anonymous):

so now you should have 0.08t = ln4.25(1.03)^t right?

OpenStudy (campbell_st):

that's correct... so the law for multiplication says add the logs \[0.08t = \log(4.25) + \log(1.03)^t\] now the law for powers \[0.08t = \ln(4.25) + t \ln(1.03)\] just collect the like terms and sole for t

OpenStudy (anonymous):

so when your adding you put t(ln(4.25 * 1.03)) right?

OpenStudy (campbell_st):

collect like terms 0.08t - t(ln(1.03)) = ln(4.25) factor the left t(0.08 - ln(1.03) ) = ln(4.25) now solve for t

OpenStudy (anonymous):

oh ok i get it thnx....i see what i did wrong